# Spatiotemporal patterns in the active cyclic Potts model

### v1

@@ Abstract
The nonequilibrium dynamics of a cycling three state Potts model is studied on a square lattice using Monte Carlo simulations and continuum theory. This model is relevant to chemical reactions on a catalytic surface and to molecular transport across a membrane. Several characteristic modes are formed depending on the flipping energies between successive states and the contact energies between neighboring sites. Under cyclic symmetry conditions, cycling homogeneous phases and spiral waves form at low and high flipping energies, respectively. In the intermediate flipping energy regime, these two modes coexist temporally in small systems and/or at low contact energies. Under asymmetric conditions, we observed small biphasic domains exhibiting amoeba-like locomotion and temporal coexistence of spiral waves and a dominant non-cyclic one-state phase. An increase in the flipping energy between two successive states, say state 0 and state 1, while keeping the other flipping energies constant, induces the formation of the third phase (state 2), owing to the suppression of the nucleation of state 0 domains. Under asymmetric conditions regarding the contact energies, two different modes can appear depending on the initial state, due to a hysteresis phenomenon.
@@ I Introduction
Spatiotemporal patterns are widely observed in nonequilibrium conditions, like waves of target and spiral shapes in two-dimensional (2D) systems Murray (2003); Mikhailov and Showalter (2006); Mikhailov and Ertl (2009); Beta and Kruse (2017); Bailles et al. (2022); Okuzono and Ohta (2003); Qu (2006); Sugimura and Kori (2015). Cyclic dynamics of three or more states are one of the conditions to generate such waves. The cyclic states are observed in various systems, such as chemical reactions, gene expression, and ecological systems Mikhailov and Showalter (2006); Mikhailov and Ertl (2009); Kerr et al. (2002); Kelsic et al. (2015). These dynamics are often explained using deterministic continuum equations Mikhailov and Showalter (2006); Mikhailov and Ertl (2009); Szolnoki et al. (2014); Reichenbach et al. (2007, 2008) and lattice Lotka–Volterra models (also called rock–paper–scissors models) Kerr et al. (2002); Kelsic et al. (2015); Szolnoki et al. (2014); Reichenbach et al. (2007, 2008); Itoh and Tainaka (1994); Tainaka (1994); Szabó et al. (1999); Szabó and Szolnoki (2002); Johnson and Seinen (2002); Szczesny et al. (2013); Juul et al. (2013); Mir et al. (2022).
@@
The spatiotemporal patterns under noise and fluctuations are not yet fully understood. Noise is typically used to trigger an excitable wave and to generate stochastic resonance Gammaitoni et al. (1998); McDonnell and Abbott (2009). The wave propagation of subexitable media of photosensitive Belousov–Zhabotinsky reaction is enhanced by the random variation of light intensity Mikhailov and Showalter (2006); Kádár et al. (1998); Wang et al. (1999); Alonso et al. (2001). In partial differential equations, random noises are added to include mesoscale fluctuations Hildebrand and Mikhailov (1996); Reichenbach et al. (2007, 2008). In the lattice Lotka–Volterra models Kerr et al. (2002); Kelsic et al. (2015); Szolnoki et al. (2014); Reichenbach et al. (2007, 2008); Itoh and Tainaka (1994); Tainaka (1994); Szabó et al. (1999); Szabó and Szolnoki (2002); Johnson and Seinen (2002); Szczesny et al. (2013); Juul et al. (2013); Mir et al. (2022) for the predator–prey systems, noise is included as a random selection of lattice sites. Since populations increase by self-reproduction, the nucleation of species domains is not considered. In contrast, in small molecular systems, thermal fluctuations have dominant effects, and nucleation and growth are the crucial kinetic processes. Spiral and target wave patterns have been observed in chemical reactions on noble metal surfaces Ertl (2008); Brär et al. (1994); Gorodetskii et al. (1994); Barroo et al. (2020); Zeininger and et al. (2022). In a previous paper Noguchi et al. , we studied a nonequilibrium three-state Potts model under cyclic symmetry (the three states being equivalent), and reported that nucleation and growth can alter spatiotemporal patterns. We found two modes: (i) a homogeneous cycling mode (HC), where each state dominantly covers the system while changing cyclically through nucleation and growth. (ii) a spiral wave (SW) mode where the three states coexist while spiral waves are formed from the contact points of three states. These two modes can temporally coexist in small systems but not in larger systems. In this study, we examine the dynamics of the three-state active Potts model in asymmetric cycling conditions. We will show several modes and hysteresis and amoeba-like motions of small biphasic domains. A non-cyclic one-state dominant phase and temporal coexistence with the SW mode newly appear.
@@
Figure 1: (a) Three-state Potts model in a square lattice. (b) Nonequilibrium cyclic flipping of the three states driven by $h_{01}+h_{12}+h_{20}=h_{\mathrm{cyc}}\neq 0$. (c) Possible realizations with energy levels depicted: (1) Chemical reaction on a catalytic surface, (2) Molecular transport through a membrane, (3) Three-state process with $1^{-}\to 1^{+}$ excitation.
@@ II Active Cyclic Potts Model
Three possible designs are described below: chemical reaction, molecular transport, and excitation, as shown in Fig. 1(c). (1) Reaction on a catalytic surface: molecules bind and unbind to the surface, with the state $s=0$ corresponding to an empty surface site, and the other two states to a site occupied by a molecule, either in the form $s=1$ or $s=2$. The surface catalyzes the reaction from $s=1$ to $s=2$, whereas this reaction is kinetically frozen in the bulk. (2) Molecular transport between two sides of a membrane: amphiphilic molecules bind to both sides of a bilayer membrane (states $s=1$ and $2$) and switch between them (flip–flop). The molecules are transported by a difference in chemical potential between the solutions on the two sides of the membrane Miele et al. (2020); Holló et al. (2021); Noguchi (2023). (3) An excitation process: $s=1$ has a low-energy state $s=1^{-}$ and a high-energy state $s=1^{+}$ triggered by external means (e.g., photoexcitation or ATP hydrolysis). Then, if the backward transformation from $s=1^{+}$ to $s=0$ is negligibly slow, we have effectively $h_{\mathrm{cyc}}=\varepsilon_{1^{+}}-\varepsilon_{1^{-}}>0$. Experimentally, chemical waves have been observed at catalytic surfaces (H2 or CO oxidation and NO or NO2 reduction on noble metal surfaces such as palladium and ruthenium) Ertl (2008); Brär et al. (1994); Gorodetskii et al. (1994); Barroo et al. (2020); Zeininger and et al. (2022). The transfer of water molecules through a chiral liquid-crystalline monolayer induces a target wave Tabe and Yokoyama (2003). Although more than two reactions or complicated molecular interactions occur in these experimental systems, we constructed a simple minimum model to capture the essence.
For simplicity, no state exchange between neighboring sites (diffusion) is considered. The thermal energy and the lattice spacing are normalized to unity. We mainly use $J_{00}=J_{11}=J_{22}=J=2$ to induce a phase separation between different states as a symmetric condition of the contact energy. To examine the effects of asymmetric contact energies, $J_{00}$ is varied from $1.75$ to $2.8$ with keeping $J_{11}=J_{22}=2$, in Sec. V. Our lattice has $N$ sites with a side length of $\sqrt{N}$ under periodic boundary conditions. The state of a site is flipped according to a Monte Carlo (MC) method. A site is chosen at random, and then it is flipped to either of the other two states with $1/2$ probability. The new state is accepted or rejected with Metropolis probability
@@
Figure 2: Cyclic flipping in symmetric condition ($h_{01}=h_{12}=h_{20}=h$ and $J_{\alpha\alpha}=2$). (a) System of size $N=256^{2}$. Probability $p_{\mathrm{phase}}=p_{\mathrm{three}}$ of the three-phase state ($N_{\alpha}>0.02N$, $N_{1}>0.02N$, and $N_{2}>0.02N$). At $h\leq 0.82$, depending on the initial condition, both the SW and the HC modes (indicated by the vanishing of $p_{\mathrm{three}}$) can appear. At $h\geq 0.83$, only the SW mode can exist. (b) System of size $N=128^{2}$. Probability $p_{\mathrm{three}}$ of the three-phase state and probability $p_{\mathrm{one}}$ of one-phase state ($N_{s}>0.98N$ with $s\in\{0,1,2\}$). The HC and SW modes temporally coexist for $1\leq h\leq 1.15$. (c) Average difference flow $q_{\mathrm{f}}$ between forward and backward flips, showing that detailed balance is more violated in the SW mode than in the HC mode. The solid and dashed lines represent the data at $N=256^{2}$ and $128^{2}$, respectively. (d) Snapshots at $h=0.9$ and $1.6$ from bottm to top for $N=256^{2}$. The light gray, blue (dark gray), and red (medium gray) represent $s=0$, $1$, and $2$, respectively.
Figure 3: Behavior of the system for low contact energy, $J_{\alpha\alpha}=1.2$, in symmetric condition ($h_{01}=h_{12}=h_{20}=h$) at $N=256^{2}$. (a) Sequential snapshots at $h=0.25$. (b) Probability $p_{\mathrm{three}}$ of the three-phase state and probability $p_{\mathrm{one}}$ of one-phase state ($N_{s}>0.95N$ with $s\in\{0,1,2\}$). The HC and SW modes temporally coexist for $0.17\leq h\leq 0.2$.
@@ III Dynamics in Cyclic-Symmetry Conditions
First, we briefly describe the dynamics of the active cyclic Potts model in symmetric conditions, i.e., for $h_{01}=h_{12}=h_{20}=h$ and $J_{00}=J_{11}=J_{22}=J$. Detailed results at $J=2$ are given in Ref. Noguchi et al. . For large systems ($N\geq 192^{2}$), at low $h$, either the HC or SW mode appears (see their definitions in Sec. I), depending on the initial state, while only the SW mode appears at high $h$. Transitions from the HC to SW modes were observed, but the reverse transition never occurred (within accessible simulation times). In the SW mode, spiral waves propagate from the contact points of three states, and the number of each state fluctuates around $N/3$ (see the snapshots in Fig. 2(d)). For small systems ($N\leq 154^{2}$), the HC mode appears at low $h$, the SW mode appears at high $h$, and the two modes temporally coexist at medium $h$. Representing temporal coexistence by a ‘+’, we denote this mixed state by SW+HC.
@@
The ratios of these two modes can be quantified using the number densities $N_{\alpha}/N$ of the three states. For $N_{\alpha}>0.98N$, we consider that the lattice is dominantly covered by a unique state $s=\alpha$, and we call this global state ‘one’ phase (except for the condition of $J=1.2$). The fraction of time spent in this state is denoted by $p_{\mathrm{phase}}=p_{\mathrm{one}}$ (Fig. 2). In the HC mode, this ‘one’ phase global state cyclically changes from $s=\alpha$ to $[\alpha+1]$, where $[\alpha+1]=(\alpha+1)\bmod 3$. Since the transient dynamics of nucleation and growth is rapid, we identify $p_{\mathrm{one}}$ with the fraction of time spent in the HC mode in the cyclic-symmetry condition. Moreover, we consider that the system is in the ‘three’ phases global state when $N_{0}>0.02N$, $N_{1}>0.02N$, and $N_{2}>0.02N$. The fraction of time spent in this state is called $p_{\mathrm{phase}}=p_{\mathrm{three}}$ (Fig. 2). The ratio $p_{\mathrm{one}}/p_{\mathrm{three}}$ corresponds to the ratio of the times spent in the HC and SW modes in the cyclic-symmetry condition (see Figs. 2(a) and (b)).
@@
Since the nucleation barrier decreases with increasing $h$ (at fixed contact energy $J$) and the nucleation more frequently occurs in larger systems, the mean lifetime of the ‘one phase’ state was found to be roughly proportional to $\exp(-h)/N$ Noguchi et al. . Furthermore, the domains were found to grow with a velocity of $v_{\mathrm{wave}}\simeq 0.07h+0.009$ at $h\geq 0.5$. Hence, at high $h$ and/or large $N$, during the growth of a $s=[\alpha+1]$ domain within a $s=\alpha$ domain, the nucleation of a smaller $s=[\alpha+2]$ domain often starts. Because domain growth is a stochastic process, three-state contacts are thus frequently produced. When a three-state contact appears, a spiral wave forms, as the three domain boundaries associated with this contact point exhibit a rotative motion (each $s=\alpha$ phase is invaded by the adjacent $s=[\alpha+1]$ phase). Conversely, the SW mode changes into the HC mode, when the three-state contact points stochastically disappear. This disappearance occurs in small systems but not in large systems ($N\geq 192^{2}$). Similar size dependence of spiral waves was reported in excitable media Qu (2006); Sugimura and Kori (2015). Since the density of the three-state contacts linearly increases with increasing $h$ (compare two snapshots in Fig. 2(d)), the SW mode more often changes into the HC mode at smaller $h$. More detail is given in Ref. Noguchi et al. .
As the contact energy $J$ decreases, the SW mode appears at lower $h$, and the cyclic change of dominant phases in the HC mode becomes faster. Thus, the temporal coexistence of these two modes can be observed in large systems. For $N=256^{2}$, the coexistence is obtained at $J=1.2$, as shown in Fig. 3. Note that the threshold $N_{\alpha}/N=0.95$ is used for the one-phase state since the domains contain 3% of the other states at $J=1.2$ (see Fig. 3(a)). The transition between the HC and SW modes occurs at higher $h$ with increasing $J$. For $J=1.5$, the discontinuous transition is obtained at $h\simeq 3.5$ and the two modes maintain for $t\sim 10^{8}$ around the transition point owing to hysteresis. However, this transition likely becomes continuous for simulations of much longer periods.
@@
Figure 7: Probabilities to observe given fractions $N_{s}/N$ of each state in the case of asymmetric flipping energies for $J_{\alpha\alpha}=2$ and $N=128^{2}$. (a) State E2 (homogeneous $s=2$ phase) for $h_{01}=1.7$ and $h_{12}=h_{20}=1$. (b) State E2 (close to the HC mode) for $h_{01}=1.45$ and $h_{12}=h_{20}=1$. (c) Temporal coexistence of the SW and HC modes (SW+HC) for $h_{01}=1.1$ and $h_{12}=h_{20}=1$, corresponding to Fig. 4. (d) SW mode at $h_{01}=1.6$, $h_{12}=1.7$, $h_{20}=1$, corresponding to Fig. 5.
@@
Figure 10: Density of states and mode lifetimes in the case of asymmetric flipping energies as a function of $h_{01}$ for $h_{12}=h_{20}=1$, $J_{\alpha\alpha}=2$, and $N=256^{2}$. (a) Fraction of sites in each state. The gray up-pointing triangles, blue down-pointing triangles, and red circles represent the data for $s=0$, $1$, and $2$, respectively. (b) Probabilities of one-phase state and three-phase coexistence. (c) Mean lifetimes of the homogeneous phases (solid lines) and SW mode (dashed line). The dynamic modes are described at the top of the figure. (d) Sequential snapshots at $h_{01}=1.3$.

### v2

@@ Abstract
The nonequilibrium dynamics of a cycling three-state Potts model is studied on a square lattice using Monte Carlo simulations and continuum theory. This model is relevant to chemical reactions on a catalytic surface and to molecular transport across a membrane. Several characteristic modes are formed depending on the flipping energies between successive states and the contact energies between neighboring sites. Under cyclic symmetry conditions, cycling homogeneous phases and spiral waves form at low and high flipping energies, respectively. In the intermediate flipping energy regime, these two modes coexist temporally in small systems and/or at low contact energies. Under asymmetric conditions, we observed small biphasic domains exhibiting amoeba-like locomotion and temporal coexistence of spiral waves and a dominant non-cyclic one-state phase. An increase in the flipping energy between two successive states, say state 0 and state 1, while keeping the other flipping energies constant, induces the formation of the third phase (state 2), owing to the suppression of the nucleation of state 0 domains. Under asymmetric conditions regarding the contact energies, two different modes can appear depending on the initial state, due to a hysteresis phenomenon.
@@ I Introduction
Spatiotemporal patterns are widely observed in nonequilibrium conditions, like waves of target and spiral shapes in two-dimensional (2D) systems Murray (2003); Mikhailov and Showalter (2006); Mikhailov and Ertl (2009); Beta and Kruse (2017); Bailles et al. (2022); Noguchi (2024); Kuramoto (1984); Okuzono and Ohta (2003); Qu (2006); Sugimura and Kori (2015). Cyclic dynamics of three or more states are one of the conditions to generate such waves. The cyclic states are observed in various systems, such as chemical reactions, gene expression, and ecological systems Mikhailov and Showalter (2006); Mikhailov and Ertl (2009); Kerr et al. (2002); Kelsic et al. (2015). These dynamics are often explained using deterministic continuum equations Mikhailov and Showalter (2006); Mikhailov and Ertl (2009); Szolnoki et al. (2014); Reichenbach et al. (2007, 2008) and lattice Lotka–Volterra models (also called rock–paper–scissors models) Kerr et al. (2002); Kelsic et al. (2015); Szolnoki et al. (2014); Reichenbach et al. (2007, 2008); Itoh and Tainaka (1994); Tainaka (1994); Szabó et al. (1999); Szabó and Szolnoki (2002); Johnson and Seinen (2002); Szczesny et al. (2013); Juul et al. (2013); Mir et al. (2022).
@@
The spatiotemporal patterns under noise and fluctuations are not yet fully understood. Noise is typically used to trigger an excitable wave and to generate stochastic resonance Gammaitoni et al. (1998); McDonnell and Abbott (2009). The wave propagation of subexitable media of photosensitive Belousov–Zhabotinsky reaction is enhanced by the random variation of light intensity Mikhailov and Showalter (2006); Kádár et al. (1998); Wang et al. (1999); Alonso et al. (2001). In partial differential equations, random noises are added to include mesoscale fluctuations Hildebrand and Mikhailov (1996); Reichenbach et al. (2007, 2008). In the lattice Lotka–Volterra models Kerr et al. (2002); Kelsic et al. (2015); Szolnoki et al. (2014); Reichenbach et al. (2007, 2008); Itoh and Tainaka (1994); Tainaka (1994); Szabó et al. (1999); Szabó and Szolnoki (2002); Johnson and Seinen (2002); Szczesny et al. (2013); Juul et al. (2013); Mir et al. (2022) for the predator–prey systems, noise is included as a random selection of lattice sites. Since populations increase by self-reproduction, the nucleation of species domains is not considered. Hence, the extinct species do not reapear (i.e., absorbing transition). In contrast, in small molecular systems, thermal fluctuations have dominant effects, and nucleation and growth are the crucial kinetic processes. Spiral and target wave patterns have been observed in chemical reactions on noble metal surfaces Ertl (2008); Brär et al. (1994); Gorodetskii et al. (1994); Barroo et al. (2020); Zeininger and et al. (2022). In a previous paper Noguchi et al. (2024), we studied a nonequilibrium three-state Potts model under cyclic symmetry (the three states being equivalent), and reported that nucleation and growth can alter spatiotemporal patterns. We found two modes: (i) a homogeneous cycling mode (HC), where each state dominantly covers the system while changing cyclically through nucleation and growth. (ii) a spiral wave (SW) mode where the three states coexist while spiral waves are formed from the contact points of three states. These two modes can temporally coexist in small systems but not in larger systems. In this study, we examine the dynamics of the three-state active Potts model in asymmetric cycling conditions. We will show several modes including hysteresis and amoeba-like motions of small biphasic domains. A non-cyclic one-state dominant phase and temporal coexistence with the SW mode newly appear.
@@
Figure 1: (a) Three-state Potts model in a square lattice. (b) Nonequilibrium cyclic flipping of the three states driven by $h_{01}+h_{12}+h_{20}=h_{\mathrm{cyc}}\neq 0$. (c) Possible realizations with energy levels depicted: (1) Chemical reaction on a catalytic surface, (2) Molecular transport through a membrane, (3) Three-state process with $1^{-}\to 1^{+}$ excitation. The wide and narrow arrows represent the dominant (forward) and secondary (backward) processes, respectively. The dashed arrows represent kinetically frozen processes.
@@ II Active Cyclic Potts Model
Three possible designs are described below: chemical reaction, molecular transport, and excitation, as shown in Fig. 1(c). (1) Reaction on a catalytic surface: molecules bind and unbind to the surface, with the state $s=0$ corresponding to an empty surface site, and the other two states to a site occupied by a molecule, either in the form $s=1$ or $s=2$. The surface catalyzes the reaction from $s=1$ to $s=2$, whereas this reaction is kinetically frozen in the bulk. (2) Molecular transport between two sides of a membrane: amphiphilic molecules bind to both sides of a bilayer membrane (states $s=1$ and $2$) and switch between them (flip–flop). The molecules are transported by a difference in chemical potential between the solutions on the two sides of the membrane Miele et al. (2020); Holló et al. (2021); Noguchi (2023). (3) An excitation process: $s=1$ has a low-energy state $s=1^{-}$ and a high-energy state $s=1^{+}$ triggered by external means (e.g., photoexcitation or ATP hydrolysis). Then, if the backward transformation from $s=1^{+}$ to $s=0$ is negligibly slow, we have effectively $h_{\mathrm{cyc}}=\varepsilon_{1^{+}}-\varepsilon_{1^{-}}>0$. Experimentally, chemical waves have been observed at catalytic surfaces (H2 or CO oxidation and NO or NO2 reduction on noble metal surfaces such as palladium and ruthenium) Ertl (2008); Brär et al. (1994); Gorodetskii et al. (1994); Barroo et al. (2020); Zeininger and et al. (2022). The transfer of water molecules through a chiral liquid-crystalline monolayer induces a target wave Tabe and Yokoyama (2003). Although multiple reactions or complicated molecular interactions occur in these experimental systems, we constructed a simple minimum model to capture the essence.
For simplicity, no state exchange between neighboring sites (diffusion) is considered. The thermal energy and the lattice spacing are normalized to unity. We mainly use $J_{00}=J_{11}=J_{22}=J=2$ to induce a phase separation between different states when using symmetric conditions for the contact energy. To examine the effects of asymmetric contact energies, $J_{00}$ is varied from $1.75$ to $2.8$ with keeping $J_{11}=J_{22}=2$, in Sec. V. Our lattice has $N$ sites with a side length of $\sqrt{N}$ under periodic boundary conditions. The state of a site is flipped according to a Monte Carlo (MC) method. A site is chosen at random, and then it is flipped to either of the other two states with $1/2$ probability. The new state is accepted or rejected with Metropolis probability
@@
Figure 2: Cyclic flipping in symmetric condition ($h_{01}=h_{12}=h_{20}=h$ and $J_{\alpha\alpha}=2$). (a) System of size $N=256^{2}$. Probability $p_{\mathrm{phase}}=p_{\mathrm{three}}$ of the three-phase state ($N_{0}>0.02N$, $N_{1}>0.02N$, and $N_{2}>0.02N$). At $h\leq 0.82$, depending on the initial condition, both the SW and the HC modes (indicated by the vanishing of $p_{\mathrm{three}}$) can appear. At $h\geq 0.83$, only the SW mode can exist. (b) System of size $N=128^{2}$. Probability $p_{\mathrm{three}}$ of the three-phase state and probability $p_{\mathrm{one}}$ of one-phase state ($N_{s}>0.98N$ with $s\in\{0,1,2\}$). The HC and SW modes temporally coexist for $1\leq h\leq 1.15$. (c) Average difference flow $q_{\mathrm{f}}$ between forward and backward flips, showing that detailed balance is more violated in the SW mode than in the HC mode. The solid and dashed lines represent the data at $N=256^{2}$ and $128^{2}$, respectively. (d) Snapshots at $h=0.9$ and $1.6$ from bottom to top for $N=256^{2}$. The light gray, blue (dark gray), and red (medium gray) represent $s=0$, $1$, and $2$, respectively.
Figure 3: Sequential snapshots at $h_{01}=h_{12}=h_{20}=h=0.25$ and $J_{\alpha\alpha}=1.2$, in symmetric condition at $N=256^{2}$.
Figure 4: Dependence of the system state on the contact energy $J$ and the flipping energy $h$ in symmetric condition ($h_{01}=h_{12}=h_{20}=h$) at $N=256^{2}$. (a) Probabilities $p_{\mathrm{phase}}=p_{\mathrm{three}}$ (blue) of the three-phase state and $p_{\mathrm{phase}}=p_{\mathrm{one}}$ (red) of the one-phase state ($N_{s}>0.95N$ with $s\in\{0,1,2\}$) at $J_{\alpha\alpha}=1.2$ as a function of $h$. The HC and SW modes temporally coexist for $0.17\leq h\leq 0.2$. (b) Dynamic phase diagram (for an observation time $t\sim 10^{8}$). The squares (light blue) and circles (gray) represent the HC and SW modes, respectively. The pluses (blue) represent the temporal coexistence of the SW and HC modes (SW+HC). The asterisks (black) represent the hysteresis region of the SW and HC modes. Either SW or HC mode appears depending on initial states.
@@ III Dynamics in Cyclic-Symmetry Conditions
First, we briefly describe the dynamics of the active cyclic Potts model in symmetric conditions, i.e., for $h_{01}=h_{12}=h_{20}=h$ and $J_{00}=J_{11}=J_{22}=J$. Detailed results at $J=2$ are given in Ref. Noguchi et al. (2024). For large systems ($N\geq 192^{2}$), at low $h$, either the HC or SW mode appears (see their definitions in Sec. I), depending on the initial state, while only the SW mode appears at high $h$. Transitions from the HC to SW modes were observed, but the reverse transition never occurred (within accessible simulation times). In the SW mode, spiral waves propagate from the contact points of three states, and the number of each state fluctuates around $N/3$ (see the snapshots in Fig. 2(d)). For small systems ($N\leq 154^{2}$), the HC mode appears at low $h$, the SW mode appears at high $h$, and the two modes temporally coexist at medium $h$. Representing temporal coexistence by a ‘+’, we denote this mixed state by SW+HC.
The ratios of these two modes can be quantified using the number densities $N_{\alpha}/N$ of the three states. For $N_{\alpha}>0.98N$, we consider that the lattice is dominantly covered by a unique state $s=\alpha$, and we call this global state ‘one’ phase, unless otherwise specified. The fraction of time spent in this state is denoted by $p_{\mathrm{phase}}=p_{\mathrm{one}}$ (Fig. 2). In the HC mode, this ‘one’ phase global state cyclically changes from $s=\alpha$ to $[\alpha+1]$, where $[\alpha+1]=(\alpha+1)\bmod 3$. Since the transient dynamics of nucleation and growth is rapid, we identify $p_{\mathrm{one}}$ with the fraction of time spent in the HC mode in the cyclic-symmetry condition. Moreover, we consider that the system is in the ‘three’ phases global state when $N_{0}>0.02N$, $N_{1}>0.02N$, and $N_{2}>0.02N$. The fraction of time spent in this state is called $p_{\mathrm{phase}}=p_{\mathrm{three}}$ (Fig. 2). The ratio $p_{\mathrm{one}}/p_{\mathrm{three}}$ corresponds to the ratio of the times spent in the HC and SW modes in the cyclic-symmetry condition (see Figs. 2(a) and (b)).
Since the nucleation barrier decreases with increasing $h$ (at fixed contact energy $J$) and the nucleation more frequently occurs in larger systems, the mean lifetime of the ‘one phase’ state was found to be roughly proportional to $\exp(-h)/N$ Noguchi et al. (2024). Furthermore, the domains were found to grow with a velocity of $v_{\mathrm{wave}}\simeq 0.07h+0.009$ at $h\geq 0.5$. Hence, at high $h$ and/or large $N$, during the growth of a $s=[\alpha+1]$ domain within a $s=\alpha$ domain, the nucleation of a smaller $s=[\alpha+2]$ domain often starts. Because domain growth is a stochastic process, three-state contacts are thus frequently produced. When a three-state contact appears, a spiral wave forms, as the three domain boundaries associated with this contact point exhibit a rotative motion (each $s=\alpha$ phase is invaded by the adjacent $s=[\alpha+1]$ phase). Conversely, the SW mode changes into the HC mode, when the three-state contact points stochastically disappear. This disappearance occurs in small systems but not in large systems ($N\geq 192^{2}$) on the simulation time scale ($t\sim 10^{8}$$10^{9}$). Similar size dependence of spiral waves was reported in excitable media Qu (2006); Sugimura and Kori (2015). Note that the continuous transition via the temporal coexistence likely recovers in the long-time limit ($t\to\infty$) even for large systems. Since the density of the three-state contacts linearly increases with increasing $h$ (compare two snapshots in Fig. 2(d)), the SW mode more often changes into the HC mode at lower $h$. With increasing system size $N$, the transition between the HC and SW modes occurs at lower $h$, and the coexistence region becomes narrower. In the large-size limit ($N\to\infty$), the HC mode does not likely appear for finite $h$ and $t\to\infty$ (the transition point $h\to 0$). More detail is given in Ref. Noguchi et al. (2024).
As the contact energy $J$ decreases, the SW mode appears at lower $h$, and the cyclic change of dominant phases in the HC mode becomes faster. Thus, the temporal coexistence of these two modes can be observed in large systems. For $N=256^{2}$, the coexistence is obtained at $J=1.2$, as shown in Figs. 3 and 4(a). Note that the threshold $N_{\alpha}/N=0.95$ is used for the one-phase state since the domains contain 3% of the other states at $J=1.2$ (see Fig. 3). The transition between the HC and SW modes occurs at higher $h$ with increasing $J$. For $J\gtrsim 1.5$, the discontinuous transition is obtained and the two modes maintain for $t\sim 10^{8}$ around the transition point owing to hysteresis (see Fig. 4(b)). However, this transition likely becomes continuous on a longer time scale. For $J=2$, the SW mode changes into the one-phase dominant phase at low $h$ ($h\lesssim 0.2$). This phase cycles as the HC mode in the long-time limit but does not cycle to the next phase on the simulation time scale, since the nucleation duration becomes exponentially long with decreasing $h$ Noguchi et al. (2024).
@@
Figure 8: Probabilities to observe given fractions $N_{s}/N$ of each state in the case of asymmetric flipping energies for $J_{\alpha\alpha}=2$ and $N=128^{2}$. (a) State E2 (homogeneous $s=2$ phase) for $h_{01}=1.7$ and $h_{12}=h_{20}=1$. (b) State E2 (close to the HC mode) for $h_{01}=1.45$ and $h_{12}=h_{20}=1$. (c) Temporal coexistence of the SW and HC modes (SW+HC) for $h_{01}=1.1$ and $h_{12}=h_{20}=1$, corresponding to Fig. 5. (d) SW mode at $h_{01}=1.6$, $h_{12}=1.7$, and $h_{20}=1$, corresponding to Fig. 6.
@@
Figure 11: Density of states and mode lifetimes in the case of asymmetric flipping energies as a function of $h_{01}$ for $h_{12}=h_{20}=1$, $J_{\alpha\alpha}=2$, and $N=256^{2}$. (a) Fraction of sites in each state. The gray up-pointing triangles, blue down-pointing triangles, and red circles represent the data for $s=0$, $1$, and $2$, respectively. (b) Probabilities of one-phase state and three-phase coexistence.with (c) Mean lifetimes of the homogeneous phases (solid lines) and SW mode (dashed line). The dynamic modes are described at the top of the figure. (d) Sequential snapshots at $h_{01}=1.3$.